100 l helium gas at 30 degrees celsius and the pressure of 2.00atm was transferred to a tank with volume of 50.0 L, what is the internal pressure of the tank if the hotness is maintained


1
Expert's answer
2021-07-05T03:23:25-0400

"V_1=100 \;l \\

p_1 = 2.0 \;atm \\

V_2=50.0 \;l \\

t= 30 \;C = constant \\

p_1V_1 = p_2V_2 \\

p_2 = \frac{p_1V_1}{V_2} \\

p_2 = \frac{2.0 \times 100}{50.0}= 4 \;atm"

Answer: 4 atm